# Blackbox Dataset Inference for LLM

### v2

@@ 6.2. Dataset Inference Accuracy
Figure 4. Comparison of positive tainted samples and negative tainted samples identified by our method for member models (top) and non-member models (bottom) under non-IID settings.
@@
Figure 5. Comparison of member samples and non-member samples identified by DPDLLM for member models (top) and non-member models (bottom) under non-IID settings.
The baseline, DPDLLM, faces problems detecting both member and non-member models, as shown in Table 6. On the Alpaca and Openhermes datasets, the accuracy drops to 42.1% and 43.9%. Even the best accuracy, achieved on Databricks, is only 71.8%. The reason, as visualized in Figure 5, is that DPDLLM presents limited accuracy when detecting the membership of individual samples. Given a member model, DPDLLM is supposed to detect all samples in $\mathcal{D}$ as members. Yet, it frequently reports more non-member samples (see top half of Figure 5). Similarly, DPDLLM can mistakenly report more samples in $\mathcal{D}$ as members, given a non-member model (see bottom half of Figure 5). Such observations are consistent with the findings by Maini et. al. (maini2024llm, ).
Table 7. Dataset inference results under the IID setting. “Non-members (1)” represent local, non-member models that share the architectures of the reference models. “Non-members (2)” are wild, non-member models with architectures different from the reference models. A value “x/y” indicates that “x” out of “y” members or non-members are correctly identified.
@@
IID: The dataset inference results under IID settings are presented in Table 7. Our method largely maintains its performance, correctly detecting all non-member models in most cases. The only exceptions occur when handling wild non-member models with Slimorca and Openhermes as the target datasets. We mistakenly report 2 and 3 out of 20 non-members as members. As shown in Figure 6, our method tends to incorrectly detect more positive tainted samples, given the two datasets under the IID setting. This problem can be alleviated by including more reference models and diversifying their architectures, as we will showcase in §6.3.
@@
Figure 6. Distribution of positive tainted samples and negative tainted samples identified by our method for non-member models under IID settings.
Figure 7. Distribution of member samples and non-member samples identified by DPDLLM for non-member models under IID settings.
6.3. Impacts of Method Designs
6.3.1. Tainted Samples
@@
Table 8. Dataset inference results without selecting tainted samples (under non-IID setting). “HF” refers to Hugging Face. A value “x/y” indicates that “x” out of “y” members or non-members are correctly identified.
Dataset Databricks Alpaca Slimorca Openhermes
Local Members 0/10 0/10 0/10 0/10
HF Members 0/10 0/10 0/10 0/10
HF Non-members 20/20 20/20 20/20 20/20
Precision (%) N/A N/A N/A N/A
Recall (%) 0.0 0.0 0.0 0.0
F1 (%) N/A N/A N/A N/A
Figure 8. Distribution of similarity difference between the suspect model and the reference models (before and after fine-tuning) on all samples from the target dataset. Given a sample, we calculate a similarity score between the suspect model’s response and the non-fine-tuned reference models’ responses. Likewise, we calculate another similarity score using the fine-tuned reference models. The difference between the two scores for all samples is used to plot the distribution.
As shown in Table 8, our method fails when all samples are used. Specifically, our method will detect every model as a non-member, regardless of the ground truth. The results are not surprising. Most samples in $\mathcal{D}$ do not impact the behaviors of the reference models. Hence, the similarity between the suspect model and the reference models on those samples doesn’t vary much before and after we fine-tune the reference models. As a demonstration,  Figure 8 shows the distribution of this similarity difference across all samples, using a member model from Hugging Face as the suspect model. Yet, our method requires a sample to stay closer to the fine-tuned reference models by a margin of $\delta^{s}$ to be considered a positive tainted sample. As a result, our method end up detecting most samples as negative tainted samples, consistently reporting the suspect model to be a non-member.
Figure 9. Impacts of reference model number and architecture diversity on our method under IID settings. “Non-members (1)” represent local, non-member models that share the architectures of the reference models. “Non-members (2)” are wild, non-member models with architectures different from the reference models.
Figure 10. Impacts of reference model number and architecture diversity on the distribution of tainted samples under IID settings.
6.3.2. Reference Models
@@
Quantity: All our evaluations so far only used one reference model each architecture. While achieving perfect accuracy under the non-IID setting, it mis-detects several non-member models under the IID setting. In this evaluation, we explore whether using more reference models improve our accuracy under the IID settings. Specifically, we increase the number of reference model to two per architecture — one fine-tuned with the parameters described in §6.1 and another using rank of matrices (4), scaling factor (16), dropout probability (0.05), bias type (none), optimizer (paged_adamw_8bit), learning rate ($5\times 10^{-5}$), batch size (8), training epoch (4), target modules (all non-linear hidden layers).
@@
Diversity: Going beyond increasing the model number, we further add two architectures, Qwen1.5-7B and glm-10b, for the reference models. To see the accumulative effects of more models and more architectures, we prepare two reference models for each architecture, following the configurations we described above.
@@
Discussion: Our evaluations above suggest that using more reference models with diversified architectures enhances the accuracy of our method. Yet, it does not mean we should endlessly increase reference models and their architectures. A clear trend unveiled in Figure 10 is that a larger family of reference models result in fewer tainted samples. When using 7 architectures with 2 models for each architecture, we often end up with fewer than 50 tainted samples in total. Decisions based on such a limited sample size are vulnerable to even minor disturbances. For instance, flipping of a small number of negative sample to positive ones due to randomness could overturn the inference result.
@@
Figure 11. Robustness of our methods against different adversarial evasion attempts. “HF” refers to Hugging Face.
6.4. Adversarial Robustness
In practice, the adversary may adopt various countermeasure to evade our dataset inference method. In this section, we evaluate the robustness of our method against several possibilities, including rephrasing responses, changing temperature, removing tainted samples and training on subsets. We focus on the member models as evasion intends to avoid the detection of them. Further, we reuse the setup discussed in §6.1 for the reference models (5 architectures, one model each architecture).
6.4.1. Rephrasing Responses
@@
As shown in 11(a), our method still detects all member models when responses are rephrased, indicating our robustness against this countermeasure. The results are attributed to our use of BERTScore, which intends to measure the semantic similarities and, thus, preserves utilities on rephrased texts.
6.4.2. Changing Temperature
Temperature is a hyperparameter that influences an LLM’s output by scaling the model’s predicted probability distribution. A higher temperature will result in lower probability, producing more creative outputs (llmtem, ). Thus, varying the temperature is expected to alter the model behaviors, which offers another evasion option to the adversaries. We experiment with this option by setting three different temperatures, including 0.0 (greedy sampling), 0.5 (half-creative sampling), and 1.0 (full-creative sampling), for the suspect model and re-perform the evaluations. All the reference models use their default temperature 1.0.
The evaluation results are illustrated in 11(b). Our method detects all the member models, given different temperatures. It demonstrates the robustness of our method against this evasion attempt. Presumably, two designs of our method contribute to this robustness. First, our use of BERTScore enables us to inspect sematic-level similarities, reducing the impacts of creativity on wording and phrasing. Second, we obtain multiple responses from the suspect model but only use the one leading to the largest similarity difference (see §5.4). This helps reduce variations introduced by the temperature.
Figure 12. Overlap of tainted samples for Llama-3.2-3B identified by our method against evasion of removing tainted samples.
6.4.3. Removing Tainted Samples
Adversaries can also attempt to remove the tainted samples from the target dataset before training their models. A key issue is whether the adversaries will be granted access to the reference models. We argue that they should not because there is no necessity for the Arbiter–who performs dataset inference–to release the reference models to third parties. If reproduction of the inference results is mandated, the Arbiter can release the reference models after membership is detected.
@@ 6.4.3. Removing Tainted Samples
11(c) presents the evaluation results, demonstrating that removing tainted samples does not diminish our detection capability. This outcome highlights the robustness of our method against such evasion attempts. The effectiveness can be attributed to the discrepancy between the reference models used by the Arbiter and those used by the adversary, which results in minimal overlap among the identified tainted samples. Further details on the overlap of tainted samples are illustrated in Figure 12. It reveals that the evasion strategy of removing tainted samples exhibits a dual effect. On one hand, it has a seemingly beneficial impact by eliminating some positive tainted samples — for example, in the Alpaca dataset, up to 29 overlapping positive tainted samples were removed. On the other hand, it inadvertently removes negative tainted samples as well, with up to 7 such overlaps also observed in the Alpaca dataset. These opposing effects partially cancel each other out. Moreover, since the original number of positive samples significantly exceeds the number of negative samples, the resulting ratio after removal remains insufficient to alter the decision-making outcome of our method. Therefore, removing tainted samples based on the adversary’s reference models does not substantially reduce the chance that suspect models will still encounter tainted samples as recognized by the Arbiter during fine-tuning. In conclusion, concealing the reference models employed by the Arbiter offers an effective defense against evasion via tainted sample removal.
6.4.4. Training On Subsets
@@
Instead of using the complete target dataset, adversaries may only use a subset for training. This can create problems for our detection as we still identify tainted samples from the entire dataset. Yet, we suspect that subset detection is no longer suitable to be solved through dataset inference. In particular, we raise the question that whether subset detection and dataset inference always share the same answer. Take the extreme case where only one sample is used for training as an example. Conceptually, it is still a subset, but we can hardly say the dataset is used. In our opinion, inference of individual samples (MIA) is a more desired solution for subset detection.
@@
Nevertheless, we extend evaluations to understand the impact of subset training on our method. We re-train the local member models using 75% and 50% of samples randomly picked from the target dataset, and then re-run our dataset inference. The results are presented in 11(d). When 75% samples are still used in training, our method largely maintains its detection capability, missing only one member on the Slimorca dataset and one on the Openhermes dataset. When the ratio of training samples drops to 50%, our method start missing member models on all datasets. On Slimorca and Openhermes, we miss 3 out of 10. This is expected as we still consider all the data to identify tainted samples. Among those, more positive tainted samples will be excluded when a smaller subset is used for training, which turn into negative ones and eventually flip our inference result.
6.5. Efficiency
Table 9. Time cost of our dataset inference method.
Dataset Offline Online Online Ratio
Databricks 4.0h 0.25h 6.25%
Alpaca 11.5h 0.25h 2.17%
Slimorca 18.0h 0.25h 1.39%
Openhermes 18.0h 0.25h 1.39%
@@ 6.5. Efficiency
To understand the efficiency—which is important in practical use—of our method, we measure its time cost for dataset inference. Our method involves two phases: (i) an offline phase which builds the reference models and identifies the tainted samples and (ii) an online phase which tests a given suspect model. We separately measure the time cost of each phase on machines with an AMD EPYC 7513 32-Core Processor, an NVIDIA A100 (80 GB GPU memory), and 256 GB RAM. Table 9 shows the evaluation results. The offline phase often takes hours to accomplish due to the time-consuming process of training reference models. Yet, it only needs to run once for every dataset, incurring a limited impact in practice. More importantly, the online phase only needs around 25 minutes to finish, presenting a high efficiency in the more frequent operations.
7. Related Works
Membership Inference Attack on LLMs: Membership inference (MI) aims to determine whether a specific data point was used to train a given machine learning model. This technique has crucial applications, including identifying data contamination in benchmark datasets (oren2023proving, ), auditing privacy violations (steinke2024privacy, ), and detecting copyrighted material within the data used to train large language models (LLMs) (carlini2021extracting, ). While MI has been well-studied for smaller models, its effectiveness on LLMs remains an area of active research. Recent advancements, however, have introduced new methods that address this challenge, which can be grouped into the score-based membership inference attacks and reference model-based MIAs according to whether they need a reference model.

### v3

@@ VI-B Dataset Inference Accuracy
Figure 3: Comparison of positive tainted samples and negative tainted samples identified by our method for member models (top) and non-member models (bottom) under non-IID settings.
@@
Figure 4: Comparison of member samples and non-member samples identified by DPDLLM for member models (top) and non-member models (bottom) under non-IID settings.
The baseline, DPDLLM, faces problems detecting both member and non-member models, as shown in Table IV. On the Alpaca and Openhermes datasets, the accuracy drops to 42.1% and 43.9%. Even the best accuracy, achieved on Databricks, is only 71.8%. The reason, as visualized in Figure 4, is that DPDLLM presents limited accuracy when detecting the membership of individual samples. Given a member model, DPDLLM is supposed to detect all samples in $\mathcal{D}$ as members. Yet, it frequently reports more non-member samples (see top half of Figure 4). Similarly, DPDLLM can mistakenly report more samples in $\mathcal{D}$ as members, given a non-member model (see bottom half of Figure 4). Such observations are consistent with the findings by Maini et. al. [9].
TABLE V: Dataset inference results under the IID setting. “Non-members (1)” represent local, non-member models that share the architectures of the reference models. “Non-members (2)” are wild, non-member models with architectures different from the reference models. A value “x/y” indicates that “x” out of “y” members or non-members are correctly identified.
@@
IID: The dataset inference results under IID settings are presented in Table V. Our method largely maintains its performance, correctly detecting all non-member models in most cases. The only exceptions occur when handling Hugging Face non-member models with Slimorca and Openhermes as the victim datasets. We mistakenly report 2 and 3 out of 20 non-members as members. As shown in Figure 5, our method tends to incorrectly detect more positive tainted samples, given the two datasets under the IID setting. This problem can be alleviated by including more reference models and diversifying their architectures, as we will showcase in §VI-C.
@@
For MIAVLM, we focus on the Shadow Model Inference attack introduced in [13], which follows the same threat model as our work and represents the strongest approach. We observed that this approach requires a large amount of data to effectively train its MIA classifier. Specifically, it relies on response features from suspect models queried under multiple temperature values. When testing suspect models that share the same architecture as the shadow models (used to train the MIA classifier) and using responses collected at ten different temperature settings, the classifier achieved 94% accuracy. However, when the method has access to responses obtained from only two temperatures, the accuracy drops to 67%. In a black-box setting, the Arbiter typically does not have full access to or control over the temperature parameter of suspect models. Moreover, when the shadow models differ in architecture from the suspect models—which is particularly common in black-box scenarios—the accuracy degrades to random guess (50%), rendering the method ineffective. Due to these limitations, we do not include it in other evaluations.
Figure 5: Distribution of positive tainted samples and negative tainted samples identified by our method for non-member models under IID settings.
Figure 6: Distribution of member samples and non-member samples identified by DPDLLM for non-member models under IID settings.
VI-C Impacts of Method Designs
VI-C1 Tainted Samples
@@
As shown in Table XI, our method fails when all samples are used. Specifically, our method will detect every model as a non-member, regardless of the ground truth. The results are not surprising. Most samples in $\mathcal{D}$ do not impact the behaviors of the reference models. Hence, the similarity between the suspect model and the reference models on those samples doesn’t vary much before and after we fine-tune the reference models. As a demonstration,  Figure 11 shows the distribution of this similarity difference across all samples, using a member model from Hugging Face as the suspect model. Yet, our method requires a sample to stay closer to the member reference models by a margin of $\delta^{s}$ to be considered a positive tainted sample. As a result, our method end up detecting most samples as negative tainted samples, consistently reporting the suspect model to be a non-member.
Additionally, we evaluate how tainted samples affect the performance of the baseline method DPDLLM. The results in Table XII show that using tainted samples in DPDLLM does not improve its accuracy. This is because DPDLLM needs to train a binary classifier for detection. Using only tainted samples provides too little data to train a well-performing classifier, and therefore it cannot achieve good detection accuracy.
Figure 7: Impacts of reference model number and architecture diversity on our method under IID settings. “Non-members (1)” represent local, non-member models that share the architectures of the reference models. “Non-members (2)” are wild, non-member models with architectures different from the reference models.
VI-C2 Reference Models
@@
Quantity: All our evaluations so far only used one reference model each architecture. While achieving perfect accuracy under the non-IID setting, it mis-detects several non-member models under the IID setting. In this evaluation, we explore whether using more reference models improve our accuracy under the IID settings. Specifically, we increase the number of reference model to two per architecture — one fine-tuned with the parameters described in §VI-A and another using rank of matrices (4), scaling factor (16), dropout probability (0.05), bias type (none), optimizer (paged_adamw_8bit), learning rate ($5\times 10^{-5}$), batch size (8), training epoch (4), target modules (all non-linear hidden layers).
@@
Diversity: Going beyond increasing the model number, we further add two architectures, Qwen1.5-7B and glm-10b, for the reference models. To see the accumulative effects of more numbers and more architectures, we prepare two reference models for each architecture, following the configurations we described above.
@@
Hyperparameters: We also explore the impact of hyperparameters used to train the reference models. Our default hyperparameters are batch size = 8 and training epoch = 3. We include four more groups of settings with increasing batch sizes (16, 32) and decreasing training epochs (1, 2), respectively.
As shown in Figure 8, changing the batch size does not affect the accuracy of dataset inference, as it does not influence the learning effectiveness of the reference models. However, a small number of training epochs leads to slightly lower detection rates of suspect models due to insufficient learning. Under the non-IID setting, a few local members are incorrectly detected, though this does not affect detection under the IID setting. More training epochs, on the other hand, can lead to the incorrect identification of certain non-members. To understand the underlying cause, we visualize the impact of training epochs of reference models on the distribution of tainted samples in Figure 13. We observe that increasing the number of training epochs enhances the reference models’ ability to select more true positive tainted samples and filter out more false positives under non-IID settings for local members. However, it has the opposite effect under the IID setting for non-members.
(a) Impact of batch size under Non-IID settings.
(b) Impact of batch size under IID settings.
(c) Impact of training epoch under Non-IID settings.
(d) Impact of training epoch under IID settings.
Figure 8: Impact of hypeparameters for reference models on dataset inference. “HF” refers to Hugging Face. “Non-members (1)” represent local, non-member models that share the same architectures of the reference models. “Non-members (2)” are wild, non-member models with architectures different from the reference models.
Discussion: Our evaluations above suggest that using more reference models with diversified architectures enhances the accuracy of our method. Yet, it does not mean we should endlessly increase reference models and their architectures. A clear trend unveiled in Figure 12 is that a larger family of reference models result in fewer tainted samples. When using 7 architectures with 2 models for each architecture, we often end up with fewer than 50 tainted samples in total. Decisions based on such a limited sample size are vulnerable to even minor disturbances. For instance, flipping of a small number of negative sample to positive ones due to randomness could overturn the inference result. Additionally, regarding hyperparameters, the batch size has no impact on our method. However, an insufficient number of training epochs can degrade the quality of the tainted samples. Increasing the number of training epochs helps filter out false positives of tainted samples and select more true positives, and hence improves the performance of our method.
VI-C3 Dataset Size
We study how the size of the dataset available for inference affects our technique. Specifically, we randomly select 5000, 10000, and 15000 of the original dataset to perform dataset inference. In order to obtain wild member models, for each dataset, we fine-tune Hugging Face non-members listed in LABEL:tab:app:models on these subsets as Hugging Face members. Additionally, we fine-tune reference models with these subsets to select tainted samples.
As shown in Figure 14, non-member models are consistently identified correctly, whereas member models are slightly affected by the smaller size of the dataset. However, the impact is limited as our method can still achieve >>95% accuracy in most cases. With the increase of the dataset size, all member models are correctly classified. These results demonstrate the robustness of our method with respect to the size of the available dataset for inference.
@@ VI-C3 Dataset Size
Figure 9: Robustness of our methods against different adversarial evasion attempts. “HF” refers to Hugging Face.
VI-D Adversarial Robustness
In practice, the adversary may adopt various countermeasure to evade our dataset inference method. In this section, we evaluate the robustness of our method against several possibilities, including rephrasing responses, changing temperature, removing tainted samples, training on subsets, and training on multiple datasets sequentially. We focus on the member models as evasion intends to avoid the detection of them. Further, we reuse the setup discussed in §VI-A for the reference models (5 architectures, one model each architecture).
VI-D1 Rephrasing Responses
@@
As shown in 9(a), our method still detects all member models when responses are rephrased, indicating our robustness against this countermeasure. The results are attributed to our use of BERTScore, which intends to measure the semantic similarities and, thus, preserves utilities on rephrased texts.
VI-D2 Changing Temperature
Temperature is a hyperparameter that influences an LLM’s output by scaling the model’s predicted probability distribution. A higher temperature will result in lower probability, producing more creative outputs [45]. Thus, varying the temperature is expected to alter the model behaviors, which offers another evasion option to the adversaries. We experiment with this option by setting three different temperatures, including 0.0 (greedy sampling), 0.5 (half-creative sampling), and 1.0 (full-creative sampling), for the suspect model and re-perform the evaluations. All the reference models use their default temperature 1.0.
The evaluation results are illustrated in 9(b). Our method detects all the member models, given different temperatures. It demonstrates the robustness of our method against this evasion attempt. Presumably, two designs of our method contribute to this robustness. First, our use of BERTScore enables us to inspect sematic-level similarities, reducing the impacts of creativity on wording and phrasing. Second, we obtain multiple responses from the suspect model but only use the one leading to the largest similarity difference (see §V-D). This helps reduce variations introduced by the temperature.
VI-D3 Removing Tainted Samples
Adversaries can also attempt to remove the tainted samples from the victim dataset before training their models. A key issue is whether the adversaries will be granted access to the reference models. We argue that they should not because there is no necessity for the Arbiter–who performs dataset inference–to release the reference models to third parties. If reproduction of the inference results is mandated, the Arbiter can release the reference models after membership is detected.
@@ VI-D3 Removing Tainted Samples
9(c) presents the evaluation results, demonstrating that removing tainted samples does not diminish our detection capability. This outcome highlights the robustness of our method against such evasion attempts. The effectiveness can be attributed to the discrepancy between the reference models used by the Arbiter and those used by the adversary, which results in minimal overlap among the identified tainted samples. There exist many tainted samples that can be used to detect the suspect model. Using different reference models to select tainted samples leads to only a very small overlap between the resulting tainted sets. Although each set of tainted samples can independently validate the suspect model, they do not share the same tainted data. Therefore, most tainted samples selected by the arbiter remain effective for detecting the suspect model and are not affected by the adversary.
@@
Further details on the overlap of tainted samples are illustrated in Figure 15. It reveals that the evasion strategy of removing tainted samples exhibits a dual effect. On one hand, it has a seemingly beneficial impact by eliminating some positive tainted samples — for example, in the Alpaca dataset, up to 29 overlapping positive tainted samples were removed. On the other hand, it inadvertently removes negative tainted samples as well, with up to 7 such overlaps also observed in the Alpaca dataset. These opposing effects partially cancel each other out. Moreover, since the original number of positive samples significantly exceeds the number of negative samples, the resulting ratio after removal remains insufficient to alter the decision-making outcome of our method. Therefore, removing tainted samples based on the adversary’s reference models does not substantially reduce the chance that suspect models will still encounter tainted samples as recognized by the Arbiter during fine-tuning. In conclusion, concealing the reference models employed by the Arbiter offers an effective defense against evasion via tainted sample removal.
VI-D4 Training On Subsets
@@
Instead of using the complete victim dataset, adversaries may only use a subset for training. This can create problems for our detection as we still identify tainted samples from the entire dataset. Yet, we suspect that subset detection is no longer suitable to be solved through dataset inference. In particular, we raise the question that whether subset detection and dataset inference always share the same answer. Take the extreme case where only one sample is used for training as an example. Conceptually, it is still a subset, but we can hardly say the dataset is used. In our opinion, inference of individual samples (MIA) is a more desired solution for subset detection.
@@
Nevertheless, we extend evaluations to understand the impact of subset training on our method. We re-train the local member models using 75% and 50% of samples randomly picked from the victim dataset, and then re-run our dataset inference. The results are presented in 9(d). When 75% samples are still used in training, our method largely maintains its detection capability, missing only one member on the Slimorca dataset and one on the Openhermes dataset. When the ratio of training samples drops to 50%, our method start missing member models on all datasets. On Slimorca and Openhermes, we miss 3 out of 10. This is expected as we still consider all the data to identify tainted samples. Among those, more positive tainted samples will be excluded when a smaller subset is used for training, which turn into negative ones and eventually flip our inference result.
VI-D5 Training On Multiple Datasets Sequentially
Instead of training LLMs on multiple datasets simultaneously, adversaries may fine-tune the models sequentially on one dataset at a time. This approach might reduce the memorization effect on earlier datasets and potentially allow adversaries to bypass our inference method. To evaluate this scenario, we train the local models in Table X on the datasets listed in Table III in a sequential manner. Reference models are trained separately on each dataset to select the corresponding tainted samples. Each time a new dataset is used for training, we perform dataset inference on all previously used datasets to determine which ones can still be successfully detected.
As shown in Table XIII, our approach is effective at detecting the two most recently used training datasets. However, its effectiveness decreases for the third-oldest and the oldest datasets. To understand the cause of this decline, we evaluate the utility performance of two representative models on all the datasets in a sequential manner, as shown in Figure 16. We observe that the models’ utility on earlier training datasets drops significantly. For example, the BERTScore of glm-2b on the oldest dataset, Databricks, decreases from nearly 1.00 to below 0.70, denoting a degradation of over 30%. This suggests that the model may have forgotten knowledge from the dataset, making it undetectable by our method. In practice, the attack aims to fully utilize the dataset to achieve high utility; however, the above approach contradicts this goal. Therefore, as long as the member model achieves good utility performance, it can be detected by our technique.
VI-E Extension to Additional Task
We extend our method to the domain of text generation. For this evaluation, we select the popular C4 dataset [^4] https://huggingface.co/datasets/allenai/c4 provided by AllenAI on Hugging Face. Specifically, we use the English subset and randomly select 50,000 data points as $\mathcal{D}$. We then choose five wild member models and five wild non-member models, as described in Table XIV. Following our proposed dataset inference method, we use the same non-member reference models from §V to construct member reference models and identify tainted samples. The evaluation results are presented in Table XV. These results demonstrate that our method generalizes well to other tasks.
VI-F Efficiency
@@
To evaluate the practical efficiency of our method, we measure its dataset-inference time cost across two phases: (i) an offline phase that builds reference models and identifies tainted samples, and (ii) an online phase that tests a suspect model. We report each phase’s runtime on a machine equipped with an AMD EPYC 7513 32-Core CPU, an NVIDIA A100 (80 GB), and 256 GB RAM. As shown in Table XVI, the offline phase takes hours due to reference-model training but is performed only once per dataset. In contrast, the online phase completes in about 25 minutes, offering high efficiency for routine use.
VII Related Works
Membership Inference Attack on LLMs: Membership inference (MI) identifies whether a data point was used in training, with uses in detecting benchmark contamination [46], auditing privacy [19], and finding copyrighted content [14]. While well-studied in small models, MI for LLMs is emerging. Methods are mainly score-based or reference model-based.